Alyssa emptied out her change jar to buy a new binder for school. She found 147
coins consisting of dimes and nickels that totaled $11.65. Write a system of
equations to determine how many dimes, d, and how many nickels, n, Alyssa had in
her change jar? Then solve the system of equations.

Found 2 solutions by mananth, ankor@dixie-net.com:
Answer by mananth(16946)


Alyssa emptied out her change jar to buy a new binder for school. She found 147
coins consisting of dimes and nickels that totaled $11.65. Write a system of
equations to determine how many dimes, d, and how many nickels, n, Alyssa had in
her change jar? Then solve the system of equations.
She found 147coins consisting of dimes and nickels
Let the number of dimes be x
and number of nickels be y
x+y=147------------------1
that totaled $11.65.
10x+5y= 1165 (cents)------------2
Multiply (1) by 10 and subtract from (2)
10x+5y= 1165
10x+10y=1470
-5y=-305
y= 61 number of nickels
x =147-61= 86 number of dimes
CHECK
61*5+86*10 =1165 cents =$11.65





15y =




Answer by ankor@dixie-net.com(22740)

Alyssa emptied out her change jar to buy a new binder for school.
She found 147 coins consisting of dimes and nickels that totaled $11.65.
Write a system of equations to determine how many dimes, d, and how many nickels, n, Alyssa had in her change jar?
Then solve the system of equations.
:
let n - no. of nickels
let d = no. of dimes
Write an equation for each statement
" She found 147 coins consisting of dimes and nickels"
n + d = 147
"that totaled $11.65."
.05n + .10d = 11.65
:
Two equations, multiply the 2nd one by 10
n + d = 147
.5n + 1d = 116.5
------------------subtraction eliminates d, find n
.5n + 0 = 30.5
n = 30.5/.5
n = 61 nickels
:
find the dimes
147 - 61 = 86 dimes
:
:
Check in the 2nd original equation
.05(61) + .10(86) =
3.05 + 8.60 = 11.65